# Standard Language Ideology in AI-Generated Language

### v1

@@ Abstract
In this position paper, we explore standard language ideology in language generated by large language models (LLMs). First, we outline how standard language ideology is reflected and reinforced in LLMs. We then present a taxonomy of open problems regarding standard language ideology in AI-generated language with implications for minoritized language communities. We introduce the concept of standard AI-generated language ideology, the process by which AI-generated language regards Standard American English (SAE) as a linguistic default and reinforces a linguistic bias that SAE is the most “appropriate” language. Finally, we discuss tensions that remain, including reflecting on what desirable system behavior looks like, as well as advantages and drawbacks of generative AI tools imitating—or often not—different English language varieties. Throughout, we discuss standard language ideology as a manifestation of existing global power structures in and through AI-generated language before ending with questions to move towards alternative, more emancipatory digital futures.
@@ 1. Introduction
Since its public release in November 2022, ChatGPT has drawn over 100 million weekly active users globally. [^1] https://www.demandsage.com/chatgpt-statistics/ While the tool brings immense benefits, those benefits are not distributed equally across speakers of different language varieties. Rather, ChatGPT—and other generative AI language technologies—reflect standard language ideology, which reinforces a hierarchy between language varieties. In this paper, we illustrate how and in what ways standard language ideology is reinforced in AI-generated language and present a taxonomy of open problems regarding standard language ideology in AI-generated language.
@@ 2.1. Standard language ideology
Standard language ideology is a construct that reinforces a hierarchy between language varieties. As defined by (lippigreen1994, ), it is “a bias toward an abstracted, idealized, homogenous spoken language which is imposed from above” whose goal is the “suppression of variation of all kinds.” In other words, it is a common and false belief that some language varieties—usually those used by communities with more social prestige—are “better” or “more complex” than others. Linguistically, however, all language varieties are equally valid. The idea that certain languages are “better” than others ignores the fact that all language varieties are equally capable of expression (Hill2008, ). There is no “correct” or “incorrect” way of using the English language or any language; in fact, “standard” language is not spoken by any real community, but is an abstracted variety that can only be defined in contrast to the speech of marginalized communities.
@@
Despite this, certain language varieties have been institutionally privileged as more “standard” and viewed as more “appropriate” or “professional” than others. This privileged status is linked to the association with people in power that the language variety holds. In particular, economic globalization has cast English as a lingua franca (ELF), or the common language adopted by speakers of different languages (Michaud2015EnglishAA, ). The English language has been granted this dominant position in international business as well as other domains, giving it and its speakers a privileged position in international business communications (Crystal2003, ; Phillipson2010-nx, ; tietze2004spreading, ). As (Wilmot2017, ) notes, this “leads to portrayal of English as a neutral solution for overcoming linguistic diversity in these relationships, as exemplified by the concept of Business English as a lingua franca.” The privileged position of English is a result of the social and historical power that English speakers have had throughout history, especially as it relates to colonization across many parts of the world. The spread of English as the default has been framed as “linguistic imperialism” that threatens other languages and language varieties (Jenkins2013-kc, ; pennycook2004, ).
@@
Importantly, even within the English language, there is a range of language varieties, with certain varieties being granted more privileged “standard” positions. In the context of the United States (US), “Standard” American English (SAE) is considered the dominant language variety and reflects an abstracted collection of linguistic norms of middle-class, white men who have held disproportionate levels of power within the country (Baker-Bell2020-hf, ; Hill2008, ; Rosa_Flores_2017, ). Other language varieties have been devalued through their institutional subordination (Baker-Bell2020-hf, ; Hill2008, ; lippigreen1994, ). African American English (AAE) is one such language variety, among many others (e.g. Irish English, (moore2024, ); Indian English, (kutlu2020negative, ); Chicano English, (guerrero2014you, )). For example, using AAE has been linked to being denied housing due to “sounding Black” (Makoni2002-wb, ; Wright2023, ). Even when using ”standard” varieties such as SAE, marginalized people can be subjected to linguistic bias (samy2007critical, ). Ultimately, language, identity and power are linked and affect people’s lives in various ways (crenshaw1995critical, ).
@@
By misleading people to believe that some languages and language varieties are better than others, standard language ideology can perpetuate harmful patterns of linguistic discrimination and the oppression of speakers of “non-standard” varieties. Linguistic discrimination can often serve as a proxy for other forms of discrimination along lines of race, gender, nationality, class, and more (Drodowicz2024, ). This discrimination resulting from the promotion of language hierarchy can be subtle under perceptions of benevolence, such as encouraging others to speak more “appropriately,” or more obvious bigotry, such as associating certain ways of speaking with lacking intelligence (baugh2005linguistic, ). Relatedly, standard language expectations dictate access to social capital through means such as education, employment, or public office (Wiley1996, ). Given this, those who speak more closely to the standard varieties benefit from better access to such resources and opportunities.
@@ 2.2. Standard language ideology in large language models
We outline how language models reinforce standard language ideology and explore why LLMs perpetuate hierarchies between language varieties. In particular, we discuss how training data for LLMs overrepresents English, particularly SAE, alongside certain voices and perspectives. We then explore how SAE is treated as the default at a higher level, which is further reinforced by the demographics of those leading generative AI research and tech companies.
Research on language models preceding the release of ChatGPT has highlighted ways in which these models perform worse for certain speakers. In particular, language models perform worse for AAE on tasks including text generation, sentiment analysis, and parsing (Blodgett2018, ; field-etal-2021-survey, ; Groenwold2020, ). (thylstrupDetectingDirtToxicity2020, ) discuss how content moderation tools fail to capture the semantic richness of AAE, such as by making blanket assumptions about the complex semantics of reclaimed slurs. Beyond worse performance for minoritized language communities, language models can also advance stereotypes regarding speakers of a particular language variety (hofmann2024dialect, ; nee2022, ).
Performance discrepancies in LLMs are linked to the language data that underpins these technologies. English is the default language for data powering large LLMs, linked to LLMs relying on language data from the Internet. Internet data overwhelmingly represents English, and Internet use varies due to social factors. An estimated 60% of all language content on the Internet is in English, despite only  17% of people speaking English globally (nee2022, ). Meanwhile, 88% of languages have “exceptionally limited resources” in digital spaces (joshi-etal-2020-state, ).
Even if languages are well-represented in digital corpora, certain perspectives are over- or under- represented. For example, on Reddit, users are 67% male and 70% White, resulting in potential reinforcement of White, male perspectives ((bender2019benderrule, ), citing (barthel2016reddit, )). Other perspectives may be actively targeted, harassed, cyberbullied, and otherwise marginalized online. Research exposes growing concerns of online gender-based harassment with clear links between online spaces and misogyny, which can be further amplified for Black and brown women (ging2018, ; duggan2017online, ; rights2014violence, ). This harassment and resulting safety and mental health implications can result in decreased use of online spaces. Meanwhile, although Black people have historically been overrepresented on Twitter compared to other demographics in the general US population (brock2012, ), Black Tweets are still often considered “inappropriate” and are more often inaccurately flagged as hateful by automatic hate speech detection tools (davidson-etal-2019-racial, ). Beyond being inaccurate, this disproportionately censors Black speakers.
Research on language models treats English—particularly SAE—as the status quo, with work on other languages “often considered ‘language specific’ and thus reviewed as less important than equivalent work on English” (bender2019benderrule, ). This default of SAE partly reflects the people who hold disproportionate power in these spaces. White men are overrepresented in AI research both industry and academia (west2019discriminating, ). Tech companies leading the generative AI charge, such as Open AI, Google, Microsoft, and Meta, are largely headquartered in the United States with employees and corporate leadership that skew White and male (omara2022why, ). The demographics of NLP research are also skewed, with researchers disproportionately affiliated with North American and European institutions; researchers from Latin America, Africa, the Middle East, and Southeast Asia are particularly underrepresented (rungta-etal-2022-geographic, ). While the demographics of those developing and leading NLP research and companies do not directly result in the reinforcement of SAE at the level of language models, this lack of diversity in the technologies’ developers and managers can lead to similar thinking in ways that miss potential biases the technology can have (cowgill2020biased, ) with a default towards an (inequitable) status quo.
2.3. Standard language ideology extends to and manifests in AI-generated language
@@
Standard language ideology extends to AI-generated language and technologies, in which hierarchies of language are reinforced. While still a relatively nascent space, there is some research that explores how generative AI technologies—including ChatGPT—perform for and exhibit biases of different language varieties. (deas-etal-2023-evaluation, ) discuss how language models including GPT-3, ChatGPT, and GPT-4 have higher perplexity for AAE, interpreted as greater difficulty in understanding AAE, and have trouble producing natural-sounding and semantically accurate AAE text. (yong-etal-2023-prompting, ) find inconsistent performance for ChatGPT’s production of Singlish [^2] Singlish is an English-based creole spoken in Singapore. and code-switched text involving Southeast Asian languages, finding that even when generated text was perceived as natural (as though produced by a native speaker), it sometimes contained “semantic inaccuracies…discernible by native speakers.” (yong-etal-2023-prompting, ) illustrate that language models produce text with harmful stereotypes about speakers of AAE through “dialect prejudice.” Another study examining performance of several widely-used GPT detectors found that non-native English writing samples were consistently misclassified as AI-generated compared to native writing samples (Liang2023, ), though the strength of this effect is uncertain.333(verma2024ghostbuster, ) suggested that the non-native writing samples being shorter had a potential confounding effect and found comparatively strong performance across models on other non-native speaker writing samples.
Examining ChatGPT performance across ten English varieties globally, [^4] Language varieties included: ”Standard” American English (SAE), African American English (AAE), ”Standard” British English (SBE), Indian English, Irish English, Jamaican English, Kenyan English, Nigerian English, Scottish English, and Singaporean English. (fleisig2024, ) found that ChatGPT responses to inputs in the different English language varieties significantly reduced the occurrence of linguistic features of all varieties except SAE and British English. SAE had the least reduction in linguistic features (22.1%) followed by British English (27.8%). Meanwhile, the other eight varieties had reductions in linguistic features of over 84%, with four languages experiencing over 96% reduction (Jamaican English, Singaporean English, Scottish English and AAE). The same study also found that, compared to Standard American and British English, ChatGPT responses to minoritized varieties of English expressed more stereotyping, demeaning, and condescending content, as well as lack of comprehension. These results illustrate that “standard” English varieties, particularly SAE, are the default output in ChatGPT, and that ChatGPT perpetuates linguistic discrimination against speakers of other varieties. Taken together, this amplifies the hierarchical position of SAE in digital spaces and technologies.
@@ 3. Taxonomy of open problems regarding standard language ideology in AI-generated language
We introduce a taxonomy of problems that standard language ideology in AI-generated language technologies present with implications for different language communities globally (see Table 1). These problems of AI-generated language technologies include: the default production of “standard” language varieties reinforces “correct” ways of communicating; outputs have lower quality of service for minoritized varieties; producing minoritized varieties can result in stereotyping of languages; producing minoritized varieties can result in appropriation and/or manipulation; and preventing outputs in minoritized varieties can result in limited quality of service and erasure.
@@
Additionally, due to automation bias—a psychological tendency to over-rely on automation that can result in complacency regarding automation outputs (goddard2012automation, )—humans can have greater trust and overconfidence in AI (horowitz2024, ). It is not clear how defaulting to SAE in generative AI language outputs impacts the mental perceptions of minoritized language speakers. However, when combined with automation bias, the default use of SAE in LLM outputs has the potential to exacerbate the idea that “standard” languages are more “correct” ways of speaking.
@@
These behaviors result in speakers of minoritized varieties experiencing more difficulty in using language models than speakers of “standard” varieties, which is a form of allocational harm (crawford2017, ). Furthermore, if users of generative AI tools are aware that using a minoritized variety results in poorer performance, they may be incentivized to use a “standard” variety instead, which can reinforce the stigmatization of non-standard varieties and result in “digital code switching.” Code switching, which results from a pressure minoritized speakers can feel to conform to more standard (more “appropriate”) language, causes psychological tolls (mccluney2019, ). While additional research is needed, these psychological tolls could extend to the digital sphere.
@@
Producing minoritized varieties can result in appropriation and/or manipulation. Even if a language model correctly produces text that reflects the grammar of that minoritized variety, there may be concerns over appropriation (particularly if non-native speakers of that variety are asking for and using this text). Because speaking a non-standard variety can carry covert prestige (i.e., social value associated with use of canonically minoritized language varieties; (labov2006social, )), use of non-standard varieties by non-native speakers, particularly in the case of White speakers appropriating AAE, has been discussed as “linguistic minstrelsy” or “figurative blackface” (Bucholtz2011, ; eberhardt2015, ). Parallel concerns have been raised regarding the use of speech, music, and image generation tools to imitate people of color, sometimes called “digital blackface” (MatamorosFernndez2022, ; sobande2021, ; Holliday2023, ; feffer2023deepdrake, ). As LLMs improve at mimicking minoritized varieties, these models may increasingly contribute to similar harms.

### v2

@@ Abstract
Standard language ideology is reflected and reinforced in language generated by large language models (LLMs). We present a faceted taxonomy of open problems that illustrate how standard language ideology manifests in AI-generated language, alongside implications for minoritized language communities and society more broadly. We introduce the concept of standard AI-generated language ideology, a process through which LLMs position ”standard” languages—and particularly Standard American English (SAE)—as the linguistic default, reinforcing the perception that SAE is the most ”appropriate” language. We then discuss ongoing tensions around what constitutes desirable system behavior, as well as advantages and drawbacks of generative AI tools attempting, or refusing, to imitate different English language varieties. Rather than prescribing narrow technical fixes, we offer three recommendations for researchers, practitioners, and funders that focus on shifting structural conditions and supporting more emancipatory outcomes for diverse language communities.
@@ 1. Introduction
Tools built on large language models (LLMs), such as ChatGPT, promise immense benefits to their users; however, those benefits are not distributed equally across speakers of different language varieties. Rather, ChatGPT—and other popular generative AI language technologies—reflect standard language ideology, which reinforces a hierarchy between language varieties and affirms ”standard” varieties (particularly ”standard” American English) as the most appropriate or legitimate form of communication. In this paper, we illustrate how and in what ways standard language ideology manifests in AI-generated language and present a novel taxonomy of open problems regarding standard language ideology in AI-generated language. Although concurrent work has surveyed challenges in natural language processing (NLP) for different languages and dialects (Joshi2025, ), this is (to our knowledge) the first to directly discuss the effects of language ideology in connection with LLMs.
We first ground our taxonomy in a literature review of research on standard language ideology, linguistic bias, and linguistic discrimination in LLMs (Section 2). This process was guided by our multidisciplinary team of researchers including linguists, computer scientists, and social scientists. We then discuss our taxonomy, which categorizes key open problems in this space and highlights their societal implications (Section 4). Finally, we discuss recommendations for developing more equitable language technologies (Section 5).
The taxonomy serves two primary purposes. First, it provides a conceptual framework for analyzing how linguistic biases in AI-generated language reflect and reinforce broader systems of power. This helps researchers identify the underlying language ideologies that shape AI development. Second, it functions as a diagnostic tool by linking those structural insights to concrete patterns in AI behavior. In this way, the framework supplies a theoretical lens for understanding why these biases arise, while the diagnostic function reveals how they materialize in practice.
Through grounding in sociolinguistic theory, we connect documented harm from AI-generated language to established research on language ideology. This connection allows us to show how well-documented hierarchies in language—including beliefs around which languages are considered ”standard” or ”better” and linguistic discrimination for speakers of minoritized varieties—are systematically reproduced in generative AI. By situating these dynamics within the framework of standard language ideology, we argue that such issues are not merely technical problems to be solved through engineering fixes, but structural problems that demand critical engagement with power and ideology. Accordingly, our recommendations aim not to mitigate particular harms, but to challenge and reshape the conditions that produce them.
@@ 2.1. Standard language ideology
Standard language ideology is a construct that reinforces a hierarchy between language varieties. As defined by (lippigreen1994, ), it is an imposed “bias toward an abstracted, idealized, homogenous spoken language” whose goal is the “suppression of variation.” In other words, it is a common and false belief that some language varieties—usually those used by communities with more social prestige—are “better” or “more complex” than others. Linguistically, however, all language varieties are equally valid. The idea that certain languages are “better” than others ignores the fact that all language varieties are equally capable of expression (Hill2008, ). In fact, “standard” language is not spoken by any real community, but is an abstracted variety that can only be defined in contrast to the speech of marginalized communities.
@@
Despite this, certain language varieties have been institutionally privileged as more “standard” and viewed as more “appropriate” than others. This privileged status is linked to the association with people in power that the language variety holds. As economic globalization has cast English as a lingua franca, or the common language adopted by speakers of different languages (Michaud2015EnglishAA, ), discrimination against different dialects of English has become intertwined with questions of discrimination against people of different nationalities and ethnicities (Crystal2003, ; Phillipson2010-nx, ; tietze2004spreading, ). The privileged position of English is a result of the social and historical power that English speakers have had throughout history, especially as it relates to colonization across many parts of the world. The spread of English as the default, and the increasing imposition of its standard varieties worldwide, has been framed as “linguistic imperialism” that threatens other languages and language varieties (Jenkins2013-kc, ; pennycook2004, ).
@@
Research on linguistic discrimination details numerous cases of discrimination against speakers of non-standard varieties. In the United States, “Standard” American English (SAE) is considered the dominant language variety and reflects an abstracted collection of linguistic norms of middle-class, white men who have held disproportionate levels of power within the country (Baker-Bell2020-hf, ; Hill2008, ; Rosa_Flores_2017, ). Other language varieties have been devalued through their institutional subordination (Baker-Bell2020-hf, ; Hill2008, ; lippigreen1994, ). African American English (AAE) is one such language variety, among many others (e.g. Irish English, (moore2024, ); Indian English, (kutlu2020negative, ); Chicano English, (guerrero2014you, )). For example, using AAE has been linked to being denied housing due to “sounding Black” (Makoni2002-wb, ; Wright2023, ). Even when using ”standard” varieties such as SAE, marginalized people can be subjected to linguistic bias (samy2007critical, ). Though linguistic discrimination in English is most well-studied, similar discrimination occurs with varieties of other languages, such as Spanish (pascual, ; Romn2023, ) and German (Adler2019, ). Because language itself is linked to identity and power (crenshaw1995critical, ), language ideologies shape both social and linguistic structures (Woolard2020, ), in ways that can reinforce social identities and existing power imbalances.
@@
By misleading people to believe that some languages and language varieties are better than others, standard language ideology can perpetuate harmful patterns of linguistic discrimination and the oppression of speakers of “non-standard” varieties. Linguistic discrimination can often serve as a proxy for other forms of discrimination along lines of race, gender, nationality, or class, among others (Drodowicz2024, ). This discrimination resulting from the promotion of language hierarchy can be subtle under perceptions of benevolence, such as encouraging others to speak more “appropriately,” or more obvious bigotry, such as associating certain ways of speaking with lacking intelligence (baugh2005linguistic, ). Relatedly, standard language expectations dictate access to social capital through means such as education, employment, or public office (Wiley1996, ). Given this, those who speak more closely to the standard varieties benefit from better access to such resources and opportunities.
@@ 2.2. Standard language ideology in large language models
Standard language ideology is reflected in LLMs in two key ways: performance discrepancies between language varieties and biases or stereotypes in outputs. Research on language models preceding the release of ChatGPT has highlighted ways in which these models perform worse for speakers of certain English language varieties. Several studies examine performance discrepancies related to AAE. For exmaple, language models perform worse for AAE on tasks including text generation, sentiment analysis, and parsing (Blodgett2018, ; field-etal-2021-survey, ; Groenwold2020, ). (thylstrupDetectingDirtToxicity2020, ) discuss how content moderation tools fail to capture the semantic richness of AAE, such as by making blanket assumptions about the complex semantics of reclaimed slurs. (Lin2024OneLM, ; ziems-etal-2022-value, ) measure performance on different NLP tasks for queries rewritten in AAE and find that tested models’ performance worsens on the rewritten queries. Furthermore, (deas-etal-2023-evaluation, ) discuss how language models including GPT-3, ChatGPT, and GPT-4 have higher perplexity for AAE, interpreted as greater difficulty in understanding AAE, and have trouble producing natural-sounding and semantically accurate AAE text. Outside of AAE, (yong-etal-2023-prompting, ) find inconsistent performance for ChatGPT’s production of Singlish [^1] Singlish is an English-based creole spoken in Singapore. and code-switched text involving Southeast Asian languages, finding that even when generated text was perceived as natural (as though produced by a native speaker), it sometimes contained “semantic inaccuracies…discernible by native speakers.”
In addition to issues of poorer performance for minoritized language communities, language models can also advance stereotypes regarding speakers of a particular language variety (hofmann2024dialect, ; nee2022, ). (finch2025findingvoiceevaluatingafrican, ) note issues in language models’ production of AAE, finding that AAE speakers prefer chatbots speaking ”mainstream” or ”standard” English varieties over AAE-speaking ones. (hofmann2024dialect, ) illustrate that language models produce text with harmful stereotypes about speakers of AAE through “dialect prejudice.’
At a higher level and illustrating the hierarchy of English language varieties in LLMs, in a study comparing eight non-standard varieties of English to Standard American and British English, (fleisig2024, ) found that language model outputs tend to default to standard varieties, particularly SAE, even in response to non-standard varieties. Responses to the eight non-standard varieties all had over 84% reduction in linguistic features from that variety, with over 96% reduction for Jamaican, Singaporean, Scottish, and African-American English. Meanwhile, outputs default to SAE orthography. Surveying native speakers of each variety to review AI-generated outputs, they also found that outputs to non-”standard” varieties consistently exhibit issues including lack of comprehension (10% worse compared to ”standard” varieties) and increased stereotyping (16% worse), condescension (12% worse), and demeaning content (22% worse). These studies illustrate that ChatGPT perpetuates linguistic discrimination against speakers of non-“standard” varieties in ways that further standard language ideology.
Performance discrepancies and stereotypes present in LLMs are linked to the language data that underpins these technologies. English is the default language for data powering large LLMs, a consequence of LLMs relying on language data from the Internet, particularly from the Common Crawl. Internet data overwhelmingly represents English, and Internet use varies due to social factors. An estimated 60% of all language content on the Internet is in English, despite only  17% of people speaking English globally (nee2022, ). Meanwhile, 88% of languages have “exceptionally limited resources” in digital spaces (joshi-etal-2020-state, ). Non-standard varieties of English are disproportionately removed from scraped Internet data when scraped pages are filtered using blocklists (dodge-etal-2021-documenting, ). Even if languages are well-represented in digital corpora, certain perspectives are over- or under- represented. For example, on Reddit, users are 67% male and 70% White, resulting in potential reinforcement of White, male perspectives ((bender2019benderrule, ), citing (barthel2016reddit, )). Other perspectives may be actively targeted, harassed, cyberbullied, and otherwise marginalized online. Research exposes growing concerns of online gender-based harassment with clear links between online spaces and misogyny, which can be further amplified for Black and brown women (ging2018, ; duggan2017online, ; rights2014violence, ). This harassment and the resulting safety and mental health implications can result in decreased use of online spaces. Meanwhile, although Black people have historically been overrepresented on Twitter compared to other demographics in the general US population (brock2012, ), Black Tweets are still often considered “inappropriate” and are more often inaccurately flagged as hateful by automatic hate speech detection tools (davidson-etal-2019-racial, ). Beyond being inaccurate, this disproportionately censors Black speakers.
Research on language models treats English—particularly SAE—as the status quo, with work on other languages “often considered ‘language specific’ and thus reviewed as less important than equivalent work on English” (bender2019benderrule, ). This default of SAE partly reflects the people who hold disproportionate power in these spaces. Leading industry developers of LLMs, such as Open AI, Google, Microsoft, and Meta, are largely headquartered in the United States, with employees and corporate leadership that skew White and male (omara2022why, ). Demographics of AI researchers in industry and academia are likewise skewed (west2019discriminating, ), with NLP researchers from Latin America, Africa, the Middle East, and Southeast Asia particularly underrepresented (rungta-etal-2022-geographic, ). While the demographics of those developing and leading NLP research and companies do not directly result in the reinforcement of SAE at the level of language models, this lack of diversity in the technologies’ developers and managers can lead to similar thinking in ways that miss potential biases the technology can have (cowgill2020biased, ) with a default towards an (inequitable) status quo.
Taken together, research illustrates the ways that standard language ideology extends to LLMs and AI-generated language, including through poorer performance and greater stereotyping for minoritized language varieties. Meanwhile, ”standard” English varieties—and particularly SAE—remains the default output.
3. Methodology
The taxonomy introduces a conceptual framework by synthesizing empirical evidence with insights from linguistic theory, sociolinguistics, and applied linguistics. It identifies and organizes a set of open problems that illustrate how standard language ideology is reproduced in generative AI outputs, alongside their broader societal implications. To identify open problems we conducted a review of large-scale studies on linguistic bias in LLMs, alongside a literature review on standard language ideology and linguistic bias. This linguistic background is essential for understanding the broader and longer-term societal implications of language technologies. While LLMs are a recent technological development, the ideological forces they reproduce have been extensively studied in prior research on language and power. Our taxonomy builds on this to surface how those same dynamics manifest in AI. After identifying relevant problems, we categorized and refined them through collaborative, interdisciplinary analysis among our multidisciplinary team. Each identified problem was linked to broader societal implications, with particular attention to harms experienced by minoritized language communities. We then iteratively refined the categories to ensure conceptual clarity and comprehensiveness. By grounding the taxonomy in both empirical studies of LLM behavior and critical sociolinguistic and AI ethics literature, this paper offers researchers and practitioners with a structured framework for understanding and addressing the reproduction of linguistic hierarchies in generative AI systems.
@@ 4. Taxonomy of open problems regarding standard language ideology in AI-generated language
We introduce a faceted taxonomy of open problems arising from reproduction of standard language ideology in AI-generated language technologies, alongside implications for different language communities globally (see Table 1). This taxonomy is structured around recurring trends identified in existing studies of linguistic bias in generative AI, with each facet linking a distinct problem to its broader societal consequences. In this section, we explore each open problem and its societal implications in detail.
@@
Additionally, due to automation bias—a psychological tendency to over-rely on automation that can result in complacency regarding automation outputs (goddard2012automation, )—humans can have greater trust and overconfidence in AI (horowitz2024, ). If LLM outputs that prioritize SAE or disparage other dialects are seen as particularly trustworthy, they could exacerbate the idea that “standard” languages are more “correct” ways of speaking.
@@
These behaviors result in speakers of minoritized varieties experiencing more difficulty in using language models than speakers of “standard” varieties, which is a form of allocational harm (crawford2017, ). Furthermore, if users of generative AI tools are aware that using a minoritized variety results in poorer performance, they may be incentivized to use a “standard” variety instead, which can reinforce the stigmatization of non-standard varieties and result in “digital code switching.” Code switching, which results from a pressure minoritized speakers can feel to conform to more “standard” language, causes psychological tolls (mccluney2019, ). While additional research is needed, these psychological tolls could extend to the digital sphere.
@@
Producing minoritized varieties can result in appropriation and/or manipulation. Even if a language model correctly produces text that reflects the grammar of that minoritized variety, there may be concerns over appropriation (particularly if non-native speakers of that variety are asking for and using this text). Because speaking a non-standard variety can carry covert prestige (i.e., social value associated with use of canonically minoritized language varieties (labov2006social, )), use of non-standard varieties by non-native speakers, particularly White speakers appropriating AAE, has been discussed as “linguistic minstrelsy” or “figurative blackface” (Bucholtz2011, ; eberhardt2015, ). Parallel concerns have been raised regarding the use of speech, music, and image generation tools to imitate people of color, sometimes called “digital blackface” (MatamorosFernndez2022, ; sobande2021, ; Holliday2023, ; feffer2023deepdrake, ). As LLMs improve at mimicking minoritized varieties, these models may increasingly contribute to similar harms.